OpenStudy (anonymous):

who can solve this prove that the function represented para metrically by the equations x= 1/sqrt{1+t^2} - ln (1+sqrt{1+t^2})/(t) , y= t/sqrt{1+t^2} satisfies the relation y=sqrt{1+(dy/dx)^2} = dy/dx

OpenStudy (anonymous):

\[x= 1/\sqrt{1+t^2} - \ln (1+\sqrt{1+t^2})/(t) , y= t/\sqrt{1+t^2} \]

OpenStudy (anonymous):

\[y \sqrt{1+(dy/dx)^2} = dy/dx\]

OpenStudy (anonymous):

hmm... my first idea would be just find bouth differentials dx, dy, and later just find their ratio dy/dx and look if its equal to ysqrt(1+(dy/dx)2 )

OpenStudy (anonymous):

ok i'll try it